OpenStudy (anonymous):

whats the probality of getting a number less than 5 and a head when rolling a die and then tossing a coin???help

OpenStudy (anonymous):

can you find the probality of getting a number less than 5 when rolling a die can you find the probality of getting a headr when tossing a coin and then multiply both of them

OpenStudy (anonymous):

ummm idk

OpenStudy (anonymous):

i got .333

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

but i need it to be a fraction

OpenStudy (anonymous):

.3333 =1/3

OpenStudy (anonymous):

wasnt part of the answer choices so put 4/12

OpenStudy (anonymous):

4/12 =1/3